Flying against the wind, a jet travels 6440 mi in 7 hours. Flying with the wind, the same jet travels 3780 mi in 3 hours. What is the rate of the jet in still air and what is the rate of the wind?
Answer by ewatrrr(24785)


Hi
Let x and y represent the speed of the jet and wind respectively
Question states*** d = r*t
7x - 7y = 6440
3x + 3y = 3780 |multiplying thru with 7/3 to eliminate y
7x - 7y = 6440
7x + 7y = 8820
14x = 15260
x = 1090mph, speed of jet
y = (7*1090 - 6440)/7 |substituting into 1st Eq
y = 170
CHECKING our Answer***
3*1090 + 3*170 = 3780